### v1

# Pair-density-wave superconductivity and Anderson’s theorem in bilayer nickelates

Abstract
The recent experimental observations of high temperature superconductivity in bilayer nickelate have attracted lots of attentions. Previous studies have assumed a mirror symmetry $\mathcal{M}$ between the two layers and focused on uniform and clean superconducting states. Here, we show that breaking this mirror symmetry via an applied displacement field can stabilize a pair-density-wave (PDW) superconductor, which is similar to the Fulde–Ferrell–Larkin–Ovchinnikov (FFLO) state, but at zero magnetic field. Based on a mean-field analysis of a model of $d_{x^{2}-y^{2}}$ orbital with an effective inter-layer attraction, we demonstrate that the PDW phase is robust over a wide range of displacement field, interlayer hopping strengths, and electron fillings. Finally, we analyze disorder effects on interlayer superconductivity within the first Born approximation. Based on symmetry considerations, we show that pairing is weaken by disorders which break the mirror symmetry, even with unbroken time reversal symmetry. Our results establish bilayer nickelate as a tunable platform for realizing finite-momentum pairing and for exploring generalized disorder effects.
Introduction.— Bilayer nickelates have recently emerged as a promising platform for exploring high-$T_{c}$ superconductivity, following the report of $80$ K superconductivity in bulk La3Ni2O7 [1, 2, 3, 4]. Motivated by their structural similarity to cuprates—yet possessing a distinct electronic configuration with a multi-orbital nature—extensive theoretical [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 5, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 9, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78] and experimental efforts [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 88, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104] have been dedicated to characterizing this system. More recently, superconductivity with $T_{c}\sim 40$ K has also been realized at ambient pressure in strained thin films [105, 106, 107, 108], providing a unique opportunity to probe the system using more accessible measurement techniques, such as angle-resolved photoemission spectroscopy (ARPES) and scanning tunneling microscopy (STM), to elucidate its electronic and magnetic properties. On the theoretical side, a wide range of studies have been carried out, proposing various mechanisms for the observed superconductivity. One plausible scenario is interlayer $s_{\pm}$-wave pairing in the $d_{x^{2}-y^{2}}$ orbital, driven by strong interlayer superexchange coupling originating from the $d_{z^{2}}$ orbital via Hund’s coupling [5, 6, 7, 8, 9, 10, 11, 12, 13]. However, most existing studies assume mirror symmetry between the two layers, which guarantees the nesting condition necessary for interlayer pairing. The regime where mirror symmetry is significantly broken remains largely unexplored.
Figure 1: (a) Schematic illustration of a bilayer nickelate. Each layer forms a square lattice hosting the $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals. In the effective one-orbital description of in $d_{x^{2}-y^{2}}$ orbital, we only keep inplane hopping $t$, and effective interlayer hopping $t_{\perp}$, generated as a higher-order process in the underlying two-orbital model (dashed arrows). An external displacement field $D$ breaks mirror symmetry between the two layers. (b) The $d_{z^{2}}$ orbital is nearly half-filled, while the $d_{x^{2}-y^{2}}$ orbital is close to quarter filling. The energy levels are shown schematically; in reality, the $d_{x^{2}-y^{2}}$ orbital forms a dispersive band with average filling n=1xn=1-x, where $x$ denotes hole doping. (c) Fermi-surface mismatch between the two layers leads to Cooper pairing with finite momentum $\delta k_{F}=k_{F,t}-k_{F,b}$. For finite $t_{\perp}$, the two layers are hybridized and $Q$ is generally defined in the band basis.
Bilayer structures introduce internal degrees of freedom that, when symmetry constraints are relaxed, can host unconventional states such as the pair-density-wave (PDW) superconductor [109]. A typical realization of a PDW is the Fulde–Ferrell–Larkin–Ovchinnikov (FFLO) state, which relies on Zeeman splitting and explicitly breaks time-reversal symmetry [110, 111]. Identifying alternative mechanisms for time-reversal-symmetric PDW states remains a significant challenge, and the search for candidate materials is ongoing. In this work, we propose that bilayer nickelates provide a promising platform to realize PDW states by breaking mirror symmetry. Specifically, under the strong interlayer pairing scenario, the resulting superconducting state can acquire a finite center-of-mass momentum upon the application of an external electric field. The momentum of the PDW is determined by the Fermi-surface mismatch between the two layers. Using a mean-field analysis of an effective single-orbital model, we demonstrate the existence of a robust PDW state over a wide range of the phase space.
In addition, we investigate the interplay between interlayer pairing and disorder in bilayer nickelates. Anderson’s theorem [112] states that conventional $s$-wave superconductivity is robust against disorder that preserves time-reversal symmetry. This principle has been generalized to multiband and odd-parity superconductors in several studies [113, 114, 115, 116]. Here, we focus on interlayer spin-singlet pairing, where the pairing partners are related by the composite $\mathcal{M}\mathcal{T}$ symmetry, combining mirror reflection $\mathcal{M}$ and time-reversal symmetry $\mathcal{T}$. By employing the superconducting fitness formalism [117, 116] within the conventional Born approximation [118, 119], we show that Anderson’s theorem for interlayer pairing can be generalized to the $\mathcal{M}\mathcal{T}$-preserving channel. Moreover, we examine the effects of finite interlayer hopping $t_{\perp}$. We find that disorder breaking the mirror symmetry $\mathcal{M}$ consistently suppresses interlayer pairing, regardless of the presence of $t_{\perp}$.
Effective one-orbital model.— The bilayer nickelate can be described by the two-orbital model on square lattice as illustrated in Fig. 1(a) (see SM) [14]. We label $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals as $d_{1}$ and $d_{2}$. Electron filling is n=2xn=2-x per site (summed over spin), with $x\approx 0.5$ relevant to experiments, corresponding to orbital fillings $n_{1}\approx 0.5$ and $n_{2}\approx 1$. The resulting band structure hosts $\alpha$, $\beta$, and $\gamma$ Fermi pockets, as depicted in Fig. 2(d). It has been shown that $\alpha$ and $\beta$ pockets exhibit mixed orbital character, while the $\gamma$ pocket is mainly from the $d_{2}$ orbital [14].
We now construct an effective one-orbital model to describe the low-energy physics more efficiently. As argued in Ref. [5, 75], the $d_{2}$ orbital is likely close to a Mott insulating state. Therefore, the mobile carriers are dominated by the $d_{1}$ orbital. Hence we focus on the $\alpha$ and $\beta$ Fermi pockets and consider an effective single-orbital model,
$H_{0}^{\mathrm{eff}}$ $=$ $\sum_{k,l,\sigma}\xi(\bm{k})\,n_{k,l,\sigma}+\sum_{k,\sigma}\gamma(\bm{k})\,c_{k,t,\sigma}^{\dagger}c_{k,b,\sigma}+\text{H.c.},$ (1)
with
$\xi(\bm{k})$ $=$ $-2t_{x}(\cos k_{x}+\cos k_{y})-\mu,$ (2)
and
$\gamma(\bm{k})$ $=$ $-t_{\perp}(\cos k_{x}-\cos k_{y})^{2}.$ (3)
[…]
Figure 2: (a) Band dispersion of the effective one-orbital model in Eq. 1. We set $t=1$ and choose $t_{\perp}=0.2$ and $t_{\perp}=0.5$. The chemical potential is set to $\mu=-1.487$ and $\mu=-1.833$, respectively, to fix the average electron density of the $d_{1}$ orbital to $n_{1}=1-x$ with $x=0.5$. (b,c) Fermi surfaces for $t_{\perp}=0.2$ (red) and $t_{\perp}=0.5$ (blue). The interlayer hopping $t_{\perp}$ splits the two bands and induces a Lifshitz transition. For small $t_{\perp}$, two electron pockets appear near $\Gamma$, while increasing $t_{\perp}$ drives one of the electron pockets into a hole pocket at $M$. (d) For comparison, we show the Fermi surface of the two-orbital model in Eq. A, hosting the $\alpha$, $\beta$, and $\gamma$ pockets.
We further introduce an external electric displacement field that breaks mirror symmetry between the two layers, which is the central tuning parameter of our study. The resulting total noninteracting Hamiltonian becomes
$H_{\mathrm{eff}}$ $=$ $H^{\mathrm{eff}}_{0}+D\sum_{k}\left(n_{t,k}-n_{b,k}\right),$ (4)
where $D$ denotes the strength of the displacement field. In the following, we will use this effective model to analyze superconducting instabilities.
Mean-field theory.— We now turn to interaction effects. We consider a phenomenological attractive interaction between the two layers, analogous to the effective interaction in conventional Bardeen–Cooper–Schrieffer (BCS) theory. Specifically, we assume an interlayer interaction of the form,
$H_{\mathrm{int}}$ $=$ $\frac{J_{\perp}^{\mathrm{eff}}}{2}\sum_{i}\left[\bm{S}_{t,i}\cdot\bm{S}_{b,i}-\frac{1}{4}n_{t,i}n_{b,i}\right].$ (5)
While our aim is to demonstrate the qualitative possibility of a PDW state, a more quantitative description would require a strong-coupling approach based on detailed microscopic models, which we leave for future work. We emphasize that $J_{\perp}^{\mathrm{eff}}$ here is not simply the $J_{\perp}$ of the $d_{z^{2}}$ orbital; rather, it is treated as an effective attraction. A more rigorous analysis would need to incorporate the local moments from the $d_{z^{2}}$ orbital and interlayer repulsion [75]. Since we expect our qualitative results to depend primarily on the pairing symmetry, we employ this simplified model for illustrative purposes.
We perform a standard mean-field analysis using an ansatz for the interlayer pairing order parameter,
$\langle c_{t,i\uparrow}c_{b,i\downarrow}\rangle=\Delta_{\bm{Q}}\exp(i\bm{R}_{i}\cdot\bm{Q}),$ (6)
where $\bm{Q}$ denotes the center-of-mass momentum of the Cooper pair. The resulting mean-field Hamiltonian is given by,
$H_{\mathrm{MF}}=H_{\mathrm{eff}}-\frac{J_{\perp}^{\mathrm{eff}}}{2}\sum_{\bm{k}}\Delta_{\bm{Q}}\,c_{t,\bm{k}^{+}}^{\dagger}(i\sigma_{y})\,c_{b,-\bm{k}^{-}}^{\dagger}+\text{H.c.},$ (7)
where $\bm{k}^{\pm}=\bm{k}\pm\bm{Q}/2$, such that $(\bm{k}^{+},-\bm{k}^{-})$ forms a Cooper pair carrying a total momentum $\bm{Q}$. The chemical potential $\mu$ is adjusted to fix the total electron density at n=1xn=1-x. For each $\bm{Q}$, the pairing amplitude $\Delta_{\bm{Q}}$ is determined self-consistently from $H_{\mathrm{MF}}$. To identify the stable phase, we compute the mean-field free energy and compare it across different values of $\bm{Q}$. Since the system lacks perfect nesting and the Fermi-momentum mismatch $|k_{F,+}-k_{F,-}|$ is not uniquely defined, we perform our analysis with $\bm{Q}=(Q,0)$ along the $x$-direction, scanning all possible $Q\in[0,\pi]$. This mean-field approach allows us to determine the energetically favored state as a function of the system parameters.
Figure 3: (a) Zero-temperature mean-field phase diagram in the $(D,t_{\perp})$ plane. We fix $J_{\perp}=4$ and $x=1/2$. As $D$ increases, the system undergoes consecutive transitions from a uniform superconducting (SC) phase to a pair-density-wave (PDW) phase, and finally to a normal state. (b, c) Mean-field solutions that minimize the free energy. For small $t_{\perp}$, the optimal pairing momentum $Q$ is close to $\delta k_{F}$, the Fermi-momentum difference along the $x$-direction as illustrated in Fig. 2(b), while it deviates from this value as $t_{\perp}$ increases. The transition between the uniform SC and PDW phases is of first order.
Pair-density-wave.— Our mean-field analysis reveals that breaking mirror symmetry via an external electric displacement field stabilizes a PDW state, similar to the Fulde–Ferrell (FF) phase. The primary results are summarized in Fig. 3.
Fig. 3(a) presents the zero-temperature phase diagram in the $(D,t_{\perp})$ plane at $x=0.5$, which corresponds to the $d^{7.5}$ electron configuration. Three distinct phases are identified: (I) a uniform superconducting (SC) phase where $\Delta_{0}\neq 0$, (II) a pair-density-wave (PDW) phase characterized by $\Delta_{\bm{Q}}\neq 0$ with $\bm{Q}\neq 0$, and (III) a normal metallic phase with $\Delta_{0}=\Delta_{\bm{Q}}=0$ for all $\bm{Q}$. For a fixed $t_{\perp}$, increasing the displacement field $D$ drives a transition from the uniform SC phase to the PDW phase, with the PDW phase appearing in the intermediate regime. At sufficiently large $D$, strong mirror-symmetry breaking induces a significant energy splitting between the layers, which suppresses interlayer pairing and stabilizes a normal metallic state, even at zero temperature.
Increasing $t_{\perp}$ tends to destabilize the PDW phase; in the large-$t_{\perp}$ limit, the system favors a uniform $s^{\pm}$ pairing state instead. Fig. 3(b) shows the optimal ordering wave vector $\bm{Q}_{c}$. For small $t_{\perp}$, $\bm{Q}_{c}$ is close to $\delta k_{F}$, but it deviates as $t_{\perp}$ increases. This deviation occurs because $\delta k_{F}$ is defined in the bonding–antibonding band basis, $\delta k_{F}=k_{F,+}-k_{F,-}$, whereas the pairing occurs between electrons on different layers in the real-space basis.
Fig. 3(c) demonstrates that the transition between the uniform SC and PDW phases is first-order, while the transitions between the PDW and normal phases, or between the SC and normal phases, are continuous. Although the form factor in the interlayer hopping $t_{\perp}$ is motivated by the specific band structure of bilayer nickelates, we find that it does not qualitatively alter the phase diagram. Consequently, similar PDW phases could be realized more generally in other bilayer systems subject to an external electric field. Finally, Fig. 4 examines the dependence on electron filling, demonstrating that the PDW phase remains robust against doping.
Generalized Anderson theorem.— We have shown that a mirror-symmetry breaking term, such as the displacement field $D$, significantly alters the phase diagram. We now turn our attention to disorder that breaks mirror reflection symmetry only locally. For conventional $s$-wave superconductors, Anderson’s theorem states that the pairing is stable against disorder as long as time-reversal symmetry $\mathcal{T}$ remains unbroken. For our interlayer-paired superconductor, however, we demonstrate that disorder breaking mirror symmetry can suppress pairing even when time-reversal symmetry is preserved.
We consider a general spin-independent disorder potential defined in the two-layer basis,
$V_{\mathrm{dis}}$ $=$ $\sum_{i=0,1,2,3}V_{i}\,\tau_{i}\otimes\sigma_{0},$ (8)
where $\tau_{i}$ and $\sigma_{i}$ are Pauli matrices acting in layer and spin space, respectively. The symmetry properties of the disorder channels are determined by the antiunitary time-reversal operator $\mathcal{T}=i\sigma_{y}K$, where $K$ denotes complex conjugation, and the unitary mirror-reflection operator $\mathcal{M}=\tau_{x}$. Accordingly, the four disorder channels are classified as, (i) mirror-symmetric and time-reversal-symmetric ($V_{0}$), (ii) mirror-symmetric and interlayer time-reversal-symmetric ($V_{1}$), (iii) mirror-antisymmetric and time-reversal-antisymmetric ($V_{2}$), and (iv) mirror-antisymmetric and time-reversal-symmetric ($V_{3}$).
Figure 4: Filling dependence of the zero-temperature mean-field phase diagram, where we fix $t_{\perp}=0.1$ and $J_{\perp}=4$. The PDW phase appears over a wide range of filling $x$.
We treat disorder perturbatively within the first-order Born approximation. We first consider the case $t_{\perp}=0$, for which the two bands are degenerate. In this limit, following Ref. [116], the disorder-induced suppression of the critical temperature $T_{c}$ can be expressed in terms of the superconducting fitness function $F_{c}$ [117],
$\log\left(\frac{T_{c}}{T_{c,0}}\right)$ $=$ $-\frac{\pi}{4}\frac{\alpha_{\mathrm{dis}}}{T_{c,0}}\Gamma,$ (9)
where $T_{c}$ is linearly suppressed by the effective scattering rate $\Gamma=n_{\mathrm{imp}}V^{2}$, and $T_{c,0}$ denotes the critical temperature in the absence of disorder. At $t_{\perp}=0$, the dimensionless depairing coefficient $\alpha_{\mathrm{dis}}$ is given by,
$\alpha_{\mathrm{dis}}$ $=$ $\left\langle\mathrm{Tr}\left(\tilde{F}_{c}^{\dagger}(k)\tilde{F}_{c}(k)\right)\right\rangle_{\mathrm{FS}},$ (10)
with the fitness function defined as,
$F_{c}$ $=$ $V_{\mathrm{dis}}\Delta-\Delta V_{\mathrm{dis}}^{*}.$ (11)
Here, the tilde denotes normalization by the disorder strength and the superconducting gap. The depairing coefficient $\alpha_{\mathrm{dis}}$ quantifies the robustness of the superconducting state, where a larger $\alpha_{\mathrm{dis}}$ corresponds to a more fragile state, while $\alpha_{\mathrm{dis}}=0$ indicates complete robustness in the weak-disorder limit.
Disorder $\mathcal{M}$ $\mathcal{T}$ $t_{\perp}=0$ $t_{\perp}\neq 0$
$V_{0}$ ++ ++
$V_{1}$ ++ ++
$V_{2}$ - -
$V_{3}$ - ++
Table 1: Summary of the generalized Anderson theorem for interlayer spin-singlet pairing, $\Delta_{\perp}=\tau_{x}\otimes i\sigma_{y}$. The second and third columns list the symmetry properties under mirror reflection $\mathcal{M}$ and time-reversal symmetry $\mathcal{T}$ for the four disorder potentials. The fourth and fifth columns indicate the stability of interlayer pairing for the cases without ($t_{\perp}=0$) and with ($t_{\perp}\neq 0$) interlayer hopping, determined by the depairing coefficient $\alpha_{\mathrm{dis}}$. A ✓  denotes $\alpha_{\mathrm{dis}}=0$ (stable SC), while a ✗  denotes $\alpha_{\mathrm{dis}}\neq 0$ (fragile SC).
For $t_{\perp}=0$ and interlayer spin-singlet pairing, $\Delta_{\perp}=\tau_{x}\otimes i\sigma_{y}$, we derive that $F_{c}=0$ when the disorder potential possesses $\mathcal{MT}$ symmetry. This implies that interlayer spin-singlet superconductivity is stable under $\mathcal{M}\mathcal{T}$-preserving disorder scattering. Accordingly, Anderson’s theorem for interlayer pairing should be modified to require $\mathcal{M}\mathcal{T}$ symmetry preservation. Applying the same analysis to intralayer pairing, $\Delta_{\parallel}=\tau_{0}\otimes i\sigma_{y}$, we recover the conventional Anderson theorem, where the non-pair-breaking condition is simply the preservation of $\mathcal{T}$.
Another important question is how this generalized Anderson theorem is modified in the presence of finite interlayer hopping $t_{\perp}$. To address this, we further analyze $\alpha_{\mathrm{dis}}$ for $t_{\perp}\neq 0$. In this case, $\alpha_{\mathrm{dis}}$ is no longer determined solely by the commutation relation between the pairing gap and the disorder potential. Instead, additional contributions arise from the commutator between the normal-state Green’s function and the pairing gap. The results are summarized in Table 1.
As shown, only the $V_{2}$ disorder channel, which is time-reversal odd and corresponds to current-like disorder, is affected by finite $t_{\perp}$ and becomes pair-breaking. Consequently, the generalized Anderson theorem derived here remains applicable in realistic experimental settings, where disorder is typically time-reversal symmetric. Notably, for mirror-symmetry breaking disorder ($V_{3}$), $\alpha_{\mathrm{dis}}$ is nonzero for both $t_{\perp}=0$ and $t_{\perp}\neq 0$, though its quantitative value changes due to the non-zero commutator with the normal Green’s function. In real materials, individual defects likely break mirror symmetry; therefore, $T_{c}$ in bilayer nickelates is expected to be sensitive to disorder, and even higher $T_{c}$ values may be achievable in cleaner samples.
Conclusion.— In this work, we have shown that breaking mirror symmetry via a displacement field in bilayer nickelates provides a natural tuning knob for realizing a pair-density-wave (PDW) state. Within a mean-field framework, we demonstrated that the PDW phase is robust over a wide region of the parameter space. Furthermore, through an analysis of disorder effects using the superconducting fitness formalism, we generalized Anderson’s theorem to interlayer spin-singlet pairing, showing that mirror-breaking disorder suppresses superconductivity. Our results establish bilayer nickelates as a tunable platform for engineering finite-momentum superconductivity and exploring symmetry-protected robustness against disorder.
Note added.— During the finalization of this manuscript, we became aware of a preprint [78], which also investigated a possible Fulde–Ferrell (FF) state driven by a displacement field.
Acknowledgements.— H.O. and Y.-H.Z. are supported by a startup fund from Johns Hopkins University and the Alfred P. Sloan Foundation through a Sloan Research Fellowship (Y.-H.Z.).
Supplemental Material for “Pair-density-wave superconductivity and Anderson’s theorem in bilayer nickelates”
Hanbit Oh1,∗, and Ya-Hui Zhang1
1 William H. Miller III Department of Physics and Astronomy,
Johns Hopkins University, Baltimore, Maryland, 21218, USA
Appendix A I. Two orbital model
We start from a two-orbital tight-binding model on a bilayer square lattice , described by the following Hamiltonian:
$H_{0}$ $=$ $-t_{x}\sum_{l,\sigma}\sum_{\langle i,j\rangle}\left(d^{\dagger}_{1,i,l,\sigma}d_{1,j,l,\sigma}+\text{H.c.}\right)$
$-t_{z}\sum_{l,\sigma}\sum_{\langle i,j\rangle}\left(d^{\dagger}_{2,i,l,\sigma}d_{2,j,l,\sigma}+\text{H.c.}\right)$
$-t_{xz}\sum_{l,\sigma}\sum_{\langle i,j\rangle}\left((-1)^{s_{ij}}d^{\dagger}_{1,i,l,\sigma}d_{2,j,l,\sigma}+\text{H.c.}\right)$
$-t^{\perp}_{z}\sum_{i,\sigma}\left(d^{\dagger}_{2,i,t,\sigma}d_{2,i,b,\sigma}+\text{H.c.}\right)+\Delta\sum_{i}(n_{i,1}-n_{i,2}).$
Here, $l=t,b$ labels the layer index, and $\sigma=\uparrow,\downarrow$ denotes the spin. We denote $d_{1}$ and $d_{2}$ as the $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals, respectively. Interlayer hopping is included only for the $d_{2}$ orbital, reflecting its strong out-of-plane character, while direct interlayer hybridization of the $d_{1}$ orbital is negligible. The hopping parameters are estimated from density functional theory as $t_{x}=0.485$, $t_{z}=0.110$, $t_{xz}=0.239$, and $t^{\perp}_{z}=0.635$ [14]. The parameter $\Delta$ represents the crystal-field splitting between the two orbitals. The factor $s_{ij}=1$ (1-1) for bonds along the $x$ ($y$) direction. The corresponding bond-dependent sign factor $(-1)^{s_{ij}}$ encodes the symmetry of the $d_{x^{2}-y^{2}}$$d_{z^{2}}$ hybridization on the square lattice. On average, the electron filling is n=2xn=2-x per site (summed over spin), with $x\approx 0.5$ relevant to experiments, corresponding to orbital fillings $n_{1}\approx 0.5$ and $n_{2}\approx 1$. The resulting band structure hosts $\alpha$, $\beta$, and $\gamma$ Fermi pockets, as depicted in Fig. 2(d). Previous studies have shown that the $\alpha$ and $\beta$ pockets exhibit mixed orbital character, while the $\gamma$ pocket is predominantly from the $d_{2}$ orbital.
Appendix B II. Details on Mean-field theory
We consider the effective one orbital model, introduced in Eq. 1,
$H_{\mathrm{eff}}$ $=$ $\sum_{\bm{k},\sigma}\phi_{\bm{k},\sigma}^{\dagger}\left(\begin{array}[]{cc}\xi(\bm{k})+D&-\gamma(\bm{k})\\ \gamma_{(}\bm{k})&\xi(\bm{k})-D\end{array}\right)\phi_{\bm{k},\sigma}=\sum_{\bm{k}}\phi_{\bm{k},\sigma}^{\dagger}\mathcal{H}_{N}(\bm{k})\phi_{\bm{k},\sigma}$ (S4)
with $\phi_{\bm{k},\sigma}^{T}=(c_{t,k,\sigma},c_{b,k,\sigma})$ and
$\xi({\bm{k}})$ $=$ $-t(\cos k_{x}+\cos k_{y})-\mu,$ (S5)
$\gamma(\bm{k})$ $=$ $-t_{\perp}(\cos k_{x}-\cos k_{y})^{2},$ (S6)
and $D$ is displacement field strength. Note that we implicitly set $\mu$ to fix the electron density n=1xn=1-x.
We next incorporate the phenomenological attractive interaction as similar setup of BCS theory. Specifically interlayer density interaction $J_{\perp}>0$ is considered,
$H_{\mathrm{int}}$ $=$ $\frac{J_{\perp}}{2}\sum_{i}\left[S_{t,i}\cdot S_{b,i}-\frac{1}{4}n_{t,i}n_{b,i}\right].$ (S7)
We then employ a standard mean-field analysis with an ansatz for the interlayer pairing order parameter,
$\langle c_{t,i\uparrow}c_{b,i\downarrow}\rangle=-\langle c_{t,i\downarrow}c_{b,i\uparrow}\rangle=\Delta_{\bm{Q}}\exp(i\bm{R}_{i}\cdot\bm{Q}),$ (S8)
where $\bm{Q}$ denotes the center-of-mass momentum of the Cooper pair.
The mean-field Hamiltonian in Nambu basis $\Phi_{\bm{k}}=\left(\begin{array}[]{cc}\phi_{\bm{k}^{+},\uparrow}\phi_{-\bm{k}^{-},\downarrow}^{*}\end{array}\right)$ is
$H_{\mathrm{MF}}$ $=$ $\frac{1}{2}\sum_{\bm{k}\in B.Z.}\Phi_{\bm{k}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}(\bm{k})\Phi_{\bm{k}}+E_{0}$
with
$\mathcal{H}_{\mathrm{BdG}}$ $=$ $\left(\begin{array}[]{cc}\mathcal{H}_{N}(\bm{k}^{+})&\mathcal{H}_{\mathrm{pair}}\\ \mathcal{H}_{\mathrm{pair}}^{\dagger}&-\mathcal{H}_{N}^{T}(-\bm{k}^{-})\end{array}\right),\quad\mathcal{H}_{\mathrm{pair}}=-J_{\perp}\Delta\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right),$ (S13)
and constant energy term
$E_{0}=\frac{1}{2}\sum_{\bm{k}\in B.Z.}\mathrm{Tr}\left(\mathcal{H}_{N}(-\bm{k}^{-})\right)+NJ_{\perp}|\Delta|^{2}.$ (S14)
Here, $\bm{k}^{\pm}=\bm{k}\pm\bm{Q}/2$ is used.
For each $\bm{Q}$, the pairing amplitude $\Delta_{\bm{Q}}$ is determined self-consistently from $H_{\mathrm{MF}}$ and Eq. S8. To further identify the phase among all $\bm{Q}$, we should compute the mean-field free-energy,
$F_{\mathrm{MF}}$ $=$ $-T\log\left(\mathrm{Tr}\left(e^{-{H_{\mathrm{MF}}/T}}\right)\right)+E_{0},$ (S15)
and compare $F$ for different $\bm{Q}$. The above mean-field analysis allows us to determine the energetically favored state as a function of system parameters.
B.1 A. Mean-field phase diagram without form factor
We also present the zero-temperature mean-field phase diagram without the momentum-dependent form factor, taking $t_{\perp}(\bm{k})=t_{\perp}$. The resulting phase diagram exhibits the same qualitative features as that obtained with the form factor shown in Fig. 3(a). As illustrated in Fig. S1, the results show that PDW is more stable than without form factor case. This is related to the shape of Fermi-surface, where the nesting vector is more well-defined in the absence of the form factor.
Figure S1: (a) Zero-temperature mean-field phase diagram without a form factor, $t_{\perp}(\bm{k})=t_{\perp}$. We fix $J_{\perp}=4$ and $x=1/2$. (b,c) Fermi surfaces without (b) and with (c) the form factor at $D=0$. In the absence of the form factor, the nesting vector is relatively well defined, which stabilizes the PDW phase, compared to Fig. 3(a).
Appendix C III. Details on computing $\alpha_{\mathrm{dis}}$
In this section, we provide derivation of Eq. S16 and expression of depairing coefficient $\alpha_{\mathrm{dis}}$ based on the 1st order Born-approximation. $\alpha_{\mathrm{dis}}$ quantifies the slope of suppression of the critical temperature under disorder,
$\log\left(\frac{T_{c}}{T_{c,0}}\right)=-\frac{\pi}{4}\frac{\alpha_{\mathrm{dis}}}{T_{c,0}}\Gamma,$ (S16)
where $T_{c}$ is linearly suppressed by the effective scattering rate $\Gamma=n_{\mathrm{imp}}V^{2}$. Here $T_{c,0}$ denotes the critical temperature in the absence of disorder.
C.1 A. Setup and 1st order Born approximation
We again consider the two-band normal-state Hamiltonian,
$H_{N}=\xi(\bm{k})\tau_{0}+t_{\perp}(\bm{k})\tau_{x},$ (S17)
where $\tau_{i}$ ($i=0,1,2,3$) are Pauli matrices defined in the layer basis. The corresponding band dispersions are given by $\xi_{\pm}(\bm{k})=\xi(\bm{k})\pm t_{\perp}(\bm{k})$, which become degenerate in the absence of interlayer hopping, $t_{\perp}=0$.
Including interlayer pairing, the Bogoliubov–de Gennes (BdG) Hamiltonian can be written as
$H_{\mathrm{BdG}}=\left(\begin{array}[]{cc}H_{N}(\bm{k}^{+})&\Delta\\ \Delta^{\dagger}&-H_{N}^{T}(-\bm{k}^{-})\end{array}\right),$ (S20)
with $\Delta=\Delta_{0}\,\tau_{1}$.
We consider a general spin-independent disorder potential in the two-layer basis,
$V_{\mathrm{dis}}=\sum_{i=0,1,2,3}V_{i}\,\tau_{i}.$ (S21)
Here we restrict our analysis to the $s$-wave scattering channel, which is momentum independent.
It is convenient to transform from the layer basis to the band basis via a unitary transformation. Under this transformation, the Pauli matrices in the layer basis are mapped to those in the band basis as
$\left\{\tau_{0},\tau_{1},\tau_{2},\tau_{3}\right\}\rightarrow\left\{\tilde{\tau}_{0},\tilde{\tau}_{3},\tilde{\tau}_{2},-\tilde{\tau}_{1}\right\}.$ (S22)
We then employ the conventional first-order Born approximation (1BA), treating $V_{\mathrm{dis}}$ as a perturbation [118, 119]. The corresponding self-energy is given by
$\Sigma(i\omega_{n})=n_{\mathrm{imp}}\int_{\bm{k}}\tilde{V}_{\mathrm{dis}}\,G_{\mathrm{BdG}}(\bm{k},i\omega_{n})\,\tilde{V}_{\mathrm{dis}},$ (S23)
where $\tilde{V}_{\mathrm{dis}}$ denotes the disorder potential written in the Nambu basis, and $i\omega_{n}$ is the fermionic Matsubara frequency.
We then decompose the self-energy in Eq. S23 into contributions that renormalize the frequency and the pairing amplitude, denoted by $\Sigma_{\omega}$ and $\Sigma_{\Delta}$, respectively. These are given by
$\Sigma_{\omega}(\bm{k},i\omega_{n})=\lim_{\Delta_{0}\rightarrow 0}\int_{\bm{k}}\frac{n_{\mathrm{imp}}V^{2}}{2}\,\mathrm{Tr}\Big[{\widehat{V}}\,G(\bm{k},i\omega_{n})\,{\widehat{V}}\Big$], (S24)
and
$\frac{\Sigma_{\Delta}(\bm{k},i\omega_{n})}{\Delta_{0}}=-\lim_{\Delta_{0}\rightarrow 0}\int_{\bm{k}}\frac{n_{\mathrm{imp}}V^{2}}{2\Delta_{0}}\,\mathrm{Tr}\Big[{\widehat{V}}\,F(\bm{k},i\omega_{n})\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big],$ (S25)
where $G$ and $F$ denote the diagonal and off-diagonal blocks of the BdG Green’s function $\mathcal{G}_{\mathrm{BdG}}$, respectively. The limit $\Delta_{0}\rightarrow 0$ is taken since we focus on the critical temperature, assuming that the superconducting transition is a continuous second-order phase transition.
C.2 B. $t_{\perp}=0$ case
We start with the simple case $t_{\perp}=0$. In this limit, the two bands are degenerate with dispersion $\xi_{\bm{k}}$. The normal and anomalous Green’s functions $G$ and $F$ in the limit $\Delta_{0}\rightarrow 0$ are simplified as
$\lim_{\Delta_{0}\rightarrow 0}G(\bm{k},i\omega_{n})$ $=$ $-\frac{i\omega_{n}+\xi_{\bm{k}}}{\omega_{n}^{2}+\xi_{\bm{k}}^{2}},\quad\lim_{\Delta_{0}\rightarrow 0}\frac{F(\bm{k},i\omega_{n})}{\Delta_{0}}=-\frac{{\widehat{\Delta}}}{\omega_{n}^{2}+\xi_{\bm{k}}^{2}}.$ (S26)
Substituting these expressions into Eqs. S24 and S25, we obtain
$\Sigma_{\omega}$ $=$ $-\frac{i\omega_{n}}{|\omega_{n}|}\,\Gamma,\quad\Sigma_{\Delta}=\frac{\Delta_{0}}{|\omega_{n}|}\Gamma\left(1-\frac{\alpha_{\mathrm{dis}}}{4}\right),$
which lead to the renormalized frequency and pairing amplitude,
$\tilde{\omega}_{n}$ $=$ $\omega_{n}+\frac{\tilde{\omega}_{n}}{|\tilde{\omega}_{n}|}\Gamma,\quad\tilde{\Delta}_{0}=\Delta_{0}+\frac{\tilde{\Delta}_{0}}{|\tilde{\omega}_{n}|}\Gamma\left(1-\frac{\alpha_{\mathrm{dis}}}{4}\right).$ (S27)
Here $\alpha_{\mathrm{dis}}$ denotes the depairing coefficient, and $F_{c}$ is the superconducting fitness function defined as
$\alpha_{\mathrm{dis}}$ $=$ $\left\langle\mathrm{Tr}\left(\tilde{F}_{c}^{\dagger}(k)\tilde{F}_{c}(k)\right)\right\rangle_{\mathrm{FS}},$ (S28)
and
$F_{c}=V\Delta-\Delta V^{*}.$ (S29)
In the above derivation, we have used the identity
[…]
Considering the disorder channels ${\widehat{V}}=\tau_{i}$ defined in Eq. S21 and the interlayer pairing ${\widehat{\Delta}}=\tau_{1}$, we find that
$F_{c}=0,\quad\text{for all channels except }\tau_{3}.$ (S31)
This leads directly to the results summarized in Table 1. Note that the $\tau_{3}$ disorder channel breaks the composite $\mathcal{M}\mathcal{T}$ symmetry, while all other channels preserve it.
C.2.1 Renormalized gap equation and $T_{c}$
We now evaluate the critical temperature in the presence (absence) of disorder, denoted by $T_{c}$ ($T_{c,0}$). In the presence of impurity scattering, the gap equation is modified as
$\frac{1}{2g}=\frac{T_{c,0}}{\mathcal{V}}\sum_{\bm{k},\omega_{n}}\frac{1}{\omega_{n}^{2}+\xi_{\bm{k}}^{2}}\;\rightarrow\;\frac{\Delta_{0}}{2g}=\frac{T_{c}}{\mathcal{V}}\sum_{\bm{k},\omega_{n}}\frac{\tilde{\Delta}_{0}}{\tilde{\omega}_{n}^{2}+\xi_{\bm{k}}^{2}},$ (S32)
where $g>0$ denotes the attractive pairing interaction.
Replacing the momentum summation by an integral over energies near the Fermi surface, the gap equation can be rewritten as
$\frac{1}{2g}\simeq\pi N(0)T_{c,0}\!\!\sum_{|\omega_{n}|<\Lambda_{UV}}\!\!\frac{1}{|\omega_{n}|}\;\rightarrow\;\frac{\Delta_{0}}{2g}\simeq\pi N(0)T_{c}\!\!\sum_{|\omega_{n}|<\Lambda_{UV}}\!\!\frac{\tilde{\Delta}_{0}}{|\tilde{\omega}_{n}|},$
where $N(0)$ is the density of states at the Fermi level and $\Lambda_{UV}$ is an ultraviolet energy cutoff imposed on the Matsubara frequencies. The angular average over the Fermi surface is defined as $\langle\mathcal{O}\rangle_{\Omega}=\int_{FS}\frac{d\Omega}{4\pi}\mathcal{O}$.
Substituting the renormalized parameters $(\tilde{\omega}_{n},\tilde{\Delta}_{0})$, we obtain
$\frac{1}{2gN(0)}\simeq 2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\Bigg[\frac{1}{\omega_{n}}-\frac{\Gamma}{\omega_{n}^{2}}\alpha_{\mathrm{dis}}\Bigg],$ (S33)
where $\Gamma=n_{\mathrm{imp}}V^{2}$ is the effective scattering rate.
To perform the Matsubara summation in Eq. S33, we make use of the digamma function $\Psi(z)$ and its derivative $\Psi^{(1)}(z)\equiv d\Psi(z)/dz$, defined as
$\Psi(z)$ $=$ $-\xi-\sum_{n=0}^{\infty}\Bigg[\frac{1}{n+z}-\frac{1}{n+1}\Bigg],\quad\Psi^{(1)}(z)=\sum_{n=0}^{\infty}\frac{1}{(n+z)^{2}},$
where $\xi$ denotes the Euler constant. For large $|z|$, the asymptotic forms $\Psi(z)\simeq\log z$ and $\Psi^{(1)}(z)\simeq 1/z$ can be used.
Applying these identities yields
$2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\frac{1}{\omega_{n}}$ $=$ $-\Bigg[\Psi\!\left(\frac{1}{2}\right)-\Psi\!\left(\frac{1}{2}+\frac{\Lambda_{UV}}{2\pi T_{c}}\right)\Bigg]$
$\simeq$ $-\Bigg[\Psi\!\left(\frac{1}{2}\right)-\log\!\left(\frac{\Lambda_{UV}}{2\pi T_{c}}\right)\Bigg],$
and
$2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\frac{1}{\omega_{n}^{2}}$ $=$ $\frac{1}{2\pi T_{c}}\Bigg[\Psi^{(1)}\!\left(\frac{1}{2}\right)-\Psi^{(1)}\!\left(\frac{1}{2}+\frac{\Lambda_{UV}}{2\pi T_{c}}\right)\Bigg]$
$\simeq$ $\frac{1}{2\pi T_{c}}\Bigg[\frac{\pi^{2}}{2}-\frac{2\pi T_{c}}{\Lambda_{UV}}\Bigg],$
valid in the limit $T_{c}\ll\Lambda_{UV}$.
Finally, we obtain the disorder-suppressed critical temperature,
$\log\left(\frac{T_{c}}{T_{c,0}}\right)=-\frac{\pi}{4}\frac{\alpha_{\mathrm{dis}}}{T_{c,0}}\Gamma,$ (S34)
where the parameter $\alpha_{\mathrm{dis}}$ quantifies the suppression of the critical temperature due to disorder.
C.3 C. $t_{\perp}\neq 0$ case
For the case $t_{\perp}\neq 0$, the form of Eq. S26 is modified due to the band splitting, $\xi_{\pm}(\bm{k})=\xi(\bm{k})\pm|t_{\perp}|$. As a result, the final expression for the depairing coefficient $\alpha_{\mathrm{dis}}$ is also modified.
Following the setup introduced in Sec. C.1, we consider the normal and anomalous Green’s functions $G$ and $F$ in the limit $\Delta_{0}\rightarrow 0$. The normal Green’s function can be written as
$\lim_{\Delta_{0}\rightarrow 0}G(\bm{k},i\omega_{n})$ $=$ $(i\omega_{n}-\xi_{+})^{-1}\tilde{\tau}_{0}+(i\omega_{n}-\xi_{-})^{-1}\tilde{\tau}_{3}\equiv G_{N}(\bm{k},i\omega_{n}),$ (S35)
which is expressed in the band basis. Here we use the tilde notation for Pauli matrices to emphasize that they are defined in the band basis, as introduced in Eq. S22.
The anomalous Green’s function in the same limit is given by
$\lim_{\Delta_{0}\rightarrow 0}\frac{F(\bm{k},i\omega_{n})}{\Delta_{0}}$ $=$ $-G_{N}{\widehat{\Delta}}G_{N}^{*}=-G_{N}G_{N}^{*}{\widehat{\Delta}}$, (S36)
where we have used the fact that $G_{N},{\widehat{\Delta}}]=0$, since both operators are proportional to ${\widehat{\Delta}}=\tilde{\tau}_{3}$.
As shown in the previous section, the key quantity governing the robustness of superconductivity within Anderson’s theorem is the relative renormalization of the frequency and pairing self-energies. In particular, when the two renormalizations are identical (i.e., $\alpha_{\mathrm{dis}}=0$), disorder does not suppress the critical temperature $T_{c}$.
We now evaluate Eqs. S24 and S25 for the case $t_{\perp}\neq 0$, again considering interlayer pairing with ${\widehat{\Delta}}=\tilde{\tau}_{3}$. We find
$\lim_{\Delta_{0}\rightarrow 0}\Big[{\widehat{V}}\,G(\bm{k},i\omega_{n})\,{\widehat{V}}\Big$] $=$ $\Big[{\widehat{V}}\,G_{N}\,{\widehat{V}}\Big$] (S37)
$=$ $\Big[G_{N}{\widehat{V}}+[{\widehat{V}},G_{N}]\Big]{\widehat{V}}$,
and
$\lim_{\Delta_{0}\rightarrow 0}\frac{1}{\Delta_{0}}\Big[{\widehat{V}}\,F(\bm{k},i\omega_{n})\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big]$ $=$ $-\Big[{\widehat{V}}\,G_{N}G_{N}^{*}\,{\widehat{\Delta}}\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big]$ (S38)
$=$ $-\Big[\big([{\widehat{V}},G_{N}G_{N}^{*}]+G_{N}G_{N}^{*}{\widehat{V}}\big){\widehat{\Delta}}\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big].$
Comparing Eqs. (S37, S38) with the corresponding expressions for the $t_{\perp}=0$ case, we find that finite interlayer hopping generates an additional commutator term, ${\widehat{V}},G_{N}]$ and ${\widehat{V}},G_{N}G_{N}^{*}]$. Therefore, in order for superconductivity to remain robust against disorder (i.e., $\alpha_{\mathrm{dis}}=0$), the condition
$[{\widehat{V}},G_{N}]=0$ (S39)
must be additionally satisfied in addition to the condition $F_{c}=0$ obtained for $t_{\perp}=0$.
Since $G_{N}$ is expressed solely in terms of $\tilde{\tau}_{0}$ and $\tilde{\tau}_{3}$, the above condition is satisfied for disorder channels proportional to $\tilde{\tau}_{0}$ and $\tilde{\tau}_{3}$, which correspond to $\tau_{0}$ and $\tau_{1}$ in the layer basis. Combining this result with the condition in Eq. S31, we conclude that
$\alpha_{\mathrm{dis}}=0,\quad\text{for disorder channels }\tau_{0},\ \tau_{1}.$ (S40)